OpenStudy (anonymous):

Find a function that models the surface area S of a cube in terms of its volume V (On my webassign account i put the answer 6v^(2/3) and it says its wrong)

jimthompson5910 (jim_thompson5910):

I'm getting the same thing. V = volume s = side length SA = surface area V = s^3 s = V^(1/3) SA = 6s^2 SA = 6(V^(1/3))^2 SA = 6V^((1/3)*2) SA = 6V^(2/3)

jimthompson5910 (jim_thompson5910):

so there's probably a glitch somewhere

jimthompson5910 (jim_thompson5910):

or you could try using a cube instead of the fraction exponent

OpenStudy (anonymous):

Thank you, you were right the website is really strict and i put in 6v^(2/3) when i should have put 6V^(2/3) thanks.

jimthompson5910 (jim_thompson5910):

oh something silly like that? how lame but I'm glad you figured it out